In the latest format of KBC, there are 5 questions. Like you all know, each question has 4 options and only one option is correct. However, in the new format, you get a chance to answer all 5 questions irrespective of the fact that you answer it right or wrong. What is the probability of getting 2 or 3 questions right?


1
Expert's answer
2022-06-06T05:48:54-0400

Let "X=" the number of right answers: "X\sim Bin (n, p)."

Given "n=5, p=1\/4=0.25, q=1-p=0.75."


"P(X=2\ or X=3)=P(X=2)+P(X=3)"

"=\dbinom{5}{2}(0.25)^{2}(0.75)^{5-2}+\dbinom{5}{3}(0.25)^{3}(0.75)^{5-3}"

"=0.3515625"